# ABACUS: An Electronic Structure Analysis Package for the AI Era

### v1

@@ V.7 DFT+U
While DFT employing LDA or GGA functionals can often predict ground-state properties of various systems with reasonable accuracy, it falls short for strongly correlated materials, such as transition metal oxides and rare-earth compounds. In these cases, the simplified exchange-correlation functional does not adequately represent complex electron interactions, resulting in inaccuracies in predicted properties such as energy band gaps, magnetic moments, and orbital polarization, etc. The DFT+U method Cococcioni and De Gironcoli (2005); Anisimov et al. (1993); Anisimov, Aryasetiawan, and Lichtenstein (1997); Dudarev et al. (1998); Jiang et al. (2010) improves the description for strongly correlated electronic systems by adding a Hubbard-U term with computational expenses comparable to those of LDA or PBE functionals.
ABACUS incorporates two distinct implementations of the DFT+U methods with NAOs basis. The primary distinction between these two methodologies lies in the way to project the electron density onto localized states, which is a critical step in accurately capturing the effects of strong electron-electron interactions within certain atomic orbitals.
The general energy functional that forms the foundation of the DFT+U method within ABACUS is initiated as
$E_{\mathrm{LDA}+\mathrm{U}}[\rho]=E_{\mathrm{LDA}}[\rho]+E_{\mathrm{Hub}}\left[\left\{n_{m}^{I\sigma}\right\}\right]-E_{\mathrm{DC}}\left[\left\{n^{I\sigma}\right\}\right],$ (110)
@@
and the fully-localized-limit (FLL) Dudarev et al. (1998) double-counting scheme assumes that the on-site Coulomb interactions of localized electrons are fully accounted for by the Hubbard-U term, subtracting a mean-field average to avoid overestimating these interactions:
$E_{U}\left[\left\{n_{mm^{\prime}}^{I\sigma}\right\}\right]=\frac{U}{2}\sum_{I}\sum_{m,\sigma}\left\{n_{mm}^{I\sigma}-\sum_{m^{\prime}}n_{mm^{\prime}}^{I\sigma}n_{m^{\prime}m}^{I\sigma}\right\},$ (111)
@@
where the $U$ value represents an empirically adjusted Coulomb penalty specifically applied to localized electrons, aiming to correct the underestimation of electron-electron repulsion within standard DFT calculations. The most critical quantity in the above expression is the on-site density occupancy matrix $n_{mm^{\prime}}^{I\sigma}$, whose implementation is different among DFT software packages. In general, it can be expressed by introducing a local projection operator $\hat{P}^{\sigma}_{\alpha mm^{\prime}}=|\alpha^{I}_{m}\sigma\rangle\langle\alpha^{I}_{m}\sigma|$ with orbital $\alpha$ at the site $I$ and density matrix $\hat{\rho}$
$n^{I\sigma}_{\alpha mm^{\prime}}=\langle\alpha m\sigma|\hat{\rho}|\alpha m^{\prime}\sigma\rangle=\mathrm{Tr}(\hat{\rho}\hat{P}_{\alpha mm^{\prime}}^{I\sigma}).$ (112)
V.7.1 Dual Projection Method
The dual project method implemented in ABACUS Qu et al. (2022a) adopts the Mulliken charge projector to construct the on-site density occupancy matrix Han, Ozaki, and Yu (2006). This projector transforms the Kohn-Sham orbital space into a localized and correlated subspace by utilizing the NAO basis, such as the $d$ or $f$ orbitals. The Mulliken charge projector meets the sum rule, ensuring the conservation of total electronic charges when summing over all projected channels.
The Mulliken charge projector is defined as
$\hat{P}^{\sigma}_{I,mm^{\prime}}=$ $\frac{1}{4N_{\mathbf{k}}}\sum_{\mathbf{k}}\left(|\tilde{\phi}_{\mathbf{k},\beta m^{\prime}\sigma}\rangle\langle\phi_{\mathbf{k},\beta m\sigma}|+|\phi_{\mathbf{k},\beta m^{\prime}\sigma}\rangle\langle\tilde{\phi}_{\mathbf{k},\beta m\sigma}|\right.$ (113)
$+\left.|\tilde{\phi}_{\mathbf{k},\beta m\sigma}\rangle\langle\phi_{\mathbf{k},\beta m^{\prime}\sigma}|+|\phi_{\mathbf{k},\beta m\sigma}\rangle\langle\tilde{\phi}_{\mathbf{k},\beta m^{\prime}\sigma}|\right),$
where $N_{k}$ is the number of $k$ points in the Brillouin zone. $\phi_{\mathbf{k},\beta m\sigma}$ and $\tilde{\phi}_{\mathbf{k},\beta m\sigma}$ are the original and dual orbitals, respectively. $\beta$ groups the indices ${I,l,\zeta}$ with $l$ and $\zeta$ belonging to the correlated channel of the correlated atom $I$.
This method is also referred to as the “dual” orbitalsHan, Ozaki, and Yu (2006) approach due to its foundational reliance on the concept of dual orbitals, which are orthogonal to each other. This method introduces a set of dual orbitals mathematically derived from the original atomic orbitals, ensuring a biorthogonal relationship between them. The dual orbitals are defined in terms of the original atomic orbitals as
$\tilde{\phi}_{\mathbf{k}\mu}(\mathbf{r})=\sum_{\nu}\phi_{\mathbf{k}\nu}(\mathbf{r})S^{-1}_{\nu\mu}(\mathbf{k}),$ (114)
where $\mu$ is the index for dual orbital and $\mathbf{k}$ is wave vector. $\phi_{\mathbf{k}\nu}(\mathbf{r})$ is the original atomic orbital. $S_{\mu\nu}(\mathbf{k})$ is the overlap matrix in reciprocal space with $S^{-1}_{\nu\mu}(\mathbf{k})$ being its inverse.
The overlap matrix $S_{\mu\nu}(\mathbf{k})$ is given by
$S_{\mu\nu}(\mathbf{k})=\langle\phi_{\mathbf{k}\mu}|\phi_{\mathbf{k}\nu}\rangle=\sum_{\mathbf{R}}e^{-i\mathbf{k}\cdot\mathbf{R}}\langle\phi_{\mathbf{R}\mu}|\phi_{0\nu}\rangle,$ (115)
where $\mathbf{R}$ denotes the lattice vectors in the unit cell. $\phi_{\mathbf{R}\mu}(\mathbf{r})$ is a NAO centered on the $a$-th atom within the unit cell $\mathbf{R}$. The biorthogonality relation between the dual and original Bloch orbitals is
$\langle\phi_{q\mu}|\tilde{\phi}_{\mathbf{k}\nu}\rangle=\delta_{kq}\delta_{\mu\nu}.$
The spin-dependent on-site density occupancy matrix is defined by dual orbitals
$n^{\sigma}_{I,mm^{\prime}}=\frac{1}{4N_{k}}\sum_{\mathbf{k}}\Big{(}$ $\sum_{S\beta m,\mu}S_{\beta m,\mu}(\mathbf{k})\rho^{\sigma}_{\mu,\beta m^{\prime}}(\mathbf{k})$ (116)
$+\rho^{\sigma}_{\beta m,\mu}(\mathbf{k})S_{\mu,\beta m^{\prime}}(\mathbf{k})$
$+S_{\beta m^{\prime},\mu}(\mathbf{k})\rho^{\sigma}_{\mu,\beta m}(\mathbf{k})$
$+\rho^{\sigma}_{\beta m^{\prime},\mu}(\mathbf{k})S_{\mu,\beta m}(\mathbf{k})\Big{)},$
where the Kohn-Sham density matrix is
$\rho^{\sigma}_{\mu\nu}(\mathbf{k})=\sum_{n}f_{\sigma n}c_{\sigma n,\mu}S_{\mu\nu}(\mathbf{k})c_{\sigma n,\nu}^{*},$ (117)
and the effective single-particle potential is
$\Delta V_{\sigma I,mm^{\prime}}=\bar{U}_{I}\left(\frac{1}{2}\delta_{mm^{\prime}}-n_{\sigma I,mm^{\prime}}\right).$ (118)
where $\bar{U}_{I}=U_{I}-J_{I}$ is the effective interaction parameter on the correlated atom $I$.
The Hamiltonian related to the DFT+U method is expressed by
$\Delta V_{\text{eff},\mu\nu}^{\mathbf{k}\sigma}$ $=\langle\phi_{\mathbf{k}\mu}|\Delta\hat{V}_{\text{eff}}^{\mathbf{k}\sigma}|\phi_{\mathbf{k}\nu}\rangle$ (119)
$=\frac{1}{4}\sum_{Imm^{\prime}}\Delta V^{\sigma}_{Imm^{\prime}}\left(S_{\beta m,\nu}^{(\mathbf{k})}\delta_{\mu,\beta m^{\prime}}+S_{\mu,\beta m^{\prime}}^{(\mathbf{k})}\delta_{\beta m,\nu}+\right.$
$\left.S_{\beta m^{\prime},\nu}^{(\mathbf{k})}\delta_{\mu,\beta m}+S_{\mu,\beta m}^{(\mathbf{k})}\delta_{\beta m^{\prime},\nu}\right),$
$f_{\sigma n}$ is the occupation number, and $c_{\sigma n,\mu}$ are the Kohn-Sham eigenvectors. $N_{k}$ is the number of kpoints in the Brillouin zone.
For the sake of simplicity, only the single-spin index case of collinear spin is mentioned. The detailed expressions for the corrections to atomic forces and stresses and the non-collinear spin cases can be found in Ref. Qu et al., 2022a.
V.7.2 Full Projection Method
The dual-orbitals method is a robust and efficient DFT+U method. However, using a subset of NAOs as projection operators in the DFT+U method may lead to a significant basis set dependence for the U term. For example, large radius cutoffs for may lead to inadvertently description for localized electron characteristics. An alternative method involves the modulation of NAOs to construct localized projection operators is proposed and has been integrated in ABACUS since v3.6.
In the demarcation of atomic localized information, it is common practice to employ spherical truncation via the empirical setting of atomic radius. This approach systematically transforms the problem into one of modulating orbitals from NAOs relative to a predetermined atomic radius. Concurrently, this must conform to three imperative criteria: first, close correspondence with the original numerical orbitals must be ensured; second, the orbitals must satisfy the normalization condition; and third, sufficient smoothness at the cutoff of the atomic radius must be ensured.
We apply a direct truncation for the tail of the radial function $\chi(r)$ of original NAOs, followed by a normalization procedure. The parameter $\sigma$ controls the smoothing interval. A normalized function truncated at $r_{c}$ can be expressed in the following form
@@ V.7.2 Full Projection Method
$\alpha(r)=\left.\frac{\chi(r)g(r;\sigma)}{\langle\chi(r)g(r;\sigma)|\chi(r)g(r;\sigma)\rangle}\right|_{\frac{\partial\langle\alpha|\chi\rangle}{\partial\sigma}=0}$
$g(r;\sigma)=\left\{\begin{matrix}1-\exp\left(-\frac{(r-r_{c})^{2}}{2\sigma^{2}}\right)&r<r_{c}\\ 0&r\geq{}r_{c}\end{matrix}\right..$ (120)
Iteratively determining the value of $\sigma$ to solve for the modulation of orbitals such that the two-center integral results of the modulated $\alpha(r)$ and original $\chi(r)$ orbitals are maximized, $\sigma_{0}$ is thus ascertained to generate the target orbitals $\alpha_{0}(r)$.
The expression for the on-site density occupancy matrix is obtained through the full projection operator $\mathbf{P}$ in real spaceHan, Ozaki, and Yu (2006) using Kohn-Sham density matrix in real space $\rho^{\mathbf{R}\sigma}_{\mu\nu}$ as follows
@@
The contribution of the DFT+U method to the real-space Hamiltonian takes the form of
$V_{\mu\nu}^{U\sigma}(\mathbf{R})$ $=\frac{\partial E_{U}\left[\left\{n_{mm^{\prime}}^{I\sigma}\right\}\right]}{\partial\rho_{\mu\nu}(\mathbf{R})}$ (122)
$=\sum_{I}\sum_{mm^{\prime}}\frac{\partial E_{U}\left[\left\{n_{mm^{\prime}}^{I\sigma}\right\}\right]}{\partial n^{I\sigma}_{mm^{\prime}}}\frac{\partial n^{I\sigma}_{mm^{\prime}}}{\partial\rho_{\mu\nu}(\mathbf{R})}$
$=\sum_{I}\sum_{mm^{\prime}}\Delta V_{Imm^{\prime}}^{\sigma}\sum_{\mathbf{R},\mathbf{R}^{\prime}}\langle\phi^{0}_{\mu}|\alpha_{m}^{I,\mathbf{R}^{\prime}}\rangle\langle\alpha_{m^{\prime}}^{I,\mathbf{R}^{\prime}}|\phi^{\mathbf{R}}_{\nu}\rangle.$
The full projection DFT+U method based on modulated atomic orbitals enables a focused approach to localized charges, offering enhanced numerical stability, while the adjustable projection orbital radii provide users with the flexibility to tailor the method for various application scenarios.
Figure 28: The band gaps of transition metal monoxide MnO (a), FeO (b), CoO (c), and NiO (d) as influenced by the Hubbard U parameter, utilizing a rhombohedral unit cell with a type-II antiferromagnetic structure. “ABACUS-Full” presents results from the full projection method. For comparative analysis, PBE+U results for other computational methods are from Ref. Qu et al., 2022a.
We provide benchmarks of the DFT+U with full projection method for the MnO, CoO, FeO, and NiO systems. All calculations are utilized by ABACUS v3.8 with the optimized norm-conserving Vanderbilt (ONCV) pseudopotentials with SG15-v1.0 version, in conjunction with an advanced double-$\zeta$ plus polarization (DZP) atomic basis. This basis set incorporates a radial cutoff of 9.0 bohr for transition metal (TM) elements and 7.0 bohr for oxygen atoms. For these calculations, the Brillouin zone discretization kspacing is 0.15 $\text{bohr}^{-1}$ and the energy cutoff is 100 Ry.
Table 1: Band gaps (in eV) and atomic magnetism (in $\mu_{B}$) with format "value of band gap (value of atomic magnetic of Mn/Co/Fe/Ni)" of MnO, FeO, CoO, and NiO as a function of effective on-site Coulomb energy $\bar{U}$ (in eV). The experimental values are presented in the last row.
@@
Exp. $3.6-3.8$ ($4.58$) $2.4$ ($3.8$) $2.4$ ($3.32$) $4.0$/$4.3$ ($1.90$)
a) MnO: $3.6-3.8$ eV (Ref. Messick, Walker, and Glosser, 1972) and 4.58 $\mu_{B}$ (Ref. Cheetham and Hope, 1983)
b) CoO: $2.4$ eV (Ref. Bowen, Adler, and Auker, 1975) and 3.8 $\mu_{B}$ (Ref. Fender, Jacobson, and Wedgwood, 1968)
c) FeO: $2.4$ eV (Ref. Powell and Spicer, 1970) and 3.32 $\mu_{B}$ (Ref. Roth, 1958)
d) NiO: $4.0$ eV (Ref. Sawatzky and Allen, 1984)/$4.3$ eV (Ref. Hüfner et al., 1984) and 3.32 $\mu_{B}$ (Ref. Khan and Erickson, 1970)
Table 1 shows the band gap and atomic magnetism of TMOs with different values of $\bar{U}$. It is specified that all four TMOs manifest a rhombohedral unit cell embracing a type-II antiferromagnetic structure. Lattice parameters for MnO, FeO, CoO, and NiO are used from Ref Tran et al., 2006, namely 4.445 Å, 4.334 Å, 4.254 Å, and 4.171 Å, respectively. The antiferromagnetic (AFM) state is postulated to align along the [111] direction. We perform these calculations by Perdew-Burke-Ernzerhof functional Perdew, Burke, and Ernzerhof (1996) with a Hubbard U correction, while neglecting the spin-orbit coupling (SOC) effect. Comparative results are utilized by various DFT softwares Qu et al. (2022a), including VASP, Wien2k, and Quantum ESPRESSO. As is shown in Fig. 28, the effect of U correction in ABACUS is notably consistent with other softwares. The agreement of our results with experimental ones proves the method’s validity and its applicability in condensed matter physics and material sciences.
V.8 Real-Time Time-Dependent DFT
Density functional theory effectively predicts ground-state properties, like bond lengths, but struggles with excited states. For instance, DFT often underestimates semiconductor band gaps compared to experimental values due to insufficient treatment of electron-hole interactions and screening effects. Bagayoko (2014)

### v3

@@ IV.6 DFT+U
While local and semilocal density functional approximations (LDA and GGAs) can often predict ground-state properties of various systems with reasonable accuracy, they exhibit significant limitations when applied to strongly correlated materials, such as transition metal oxides and rare-earth compounds. In these systems, the simplified exchange-correlation functionals fail to adequately capture complex electron-electron interactions, leading to substantial inaccuracies in predicted electronic properties, including but not limited to energy band gaps, magnetic moments, and orbital polarization. The DFT+$U$ method Cococcioni and De Gironcoli (2005); Anisimov et al. (1993); Anisimov, Aryasetiawan, and Lichtenstein (1997); Dudarev et al. (1998); Jiang et al. (2010) addresses these limitations by incorporating a Hubbard-$U$ correction term, which significantly improves the description of strongly correlated electronic systems while maintaining computational costs comparable to standard LDA or GGA calculations.
The DFT+$U$ method is implemented through the following total energy functional
$E_{\mathrm{DFA}+U}[\rho]=E_{\mathrm{DFA}}[\rho]+E_{\mathrm{Hub}}-E_{\mathrm{dc}},$ (98)
@@
where $E_{\text{LDA}}$ denotes the baseline energy. In addition, the Hubbard term and double-counting terms are labeled as $E_{\mathrm{Hub}}$ and $E_{\mathrm{dc}}$, respectively. In ABACUS, we introduce the fully localized limit (FLL) implementation of DFT+$U$ method Dudarev et al. (1998), which assumes that the on-site Coulomb interactions of localized electrons are fully accounted for by the Hubbard $U$ term, while a mean-field average is subtracted to avoid overestimating these interactions. The Hubbard-corrected energy functional is given by
$E_{U}\left[\left\{n_{mm^{\prime}}^{I\sigma}\right\}\right]=\frac{U}{2}\sum_{I}\sum_{m\sigma}\Big{(}n_{mm}^{I\sigma}-\sum_{m^{\prime}}n_{mm^{\prime}}^{I\sigma}n_{m^{\prime}m}^{I\sigma}\Big{)}.$ (99)
@@
Here, the Hubbard parameter $U$ represents an empirically adjusted Coulomb penalty specifically applied to localized electrons, designed to correct the systematic underestimation of electron-electron repulsion inherent in standard DFT calculations. Additionally, the central quantity is the on-site density occupancy matrix $n_{mm^{\prime}}^{I\sigma}$ with $I$, $\sigma$, and $m$ being the atom, spin and angular momentum indices, respectively. In general, this matrix can be expressed using a local projection operator
$\hat{P}^{\sigma}_{mm^{\prime}}=|\alpha^{I\sigma}_{m}\rangle\langle\alpha^{I\sigma}_{m^{\prime}}|,$ (100)
where $\alpha^{I\sigma}_{m}$ denotes the projection orbital, and $\hat{\rho}$ represents the density matrix
$n^{I\sigma}_{mm^{\prime}}=\langle\alpha^{I\sigma}_{m}\sigma|\hat{\rho}|\alpha^{I\sigma}_{m^{\prime}}\sigma\rangle=\mathrm{Tr}(\hat{\rho}\hat{P}_{mm^{\prime}}^{I\sigma}).$ (101)
Two distinct implementations of the DFT+$U$ method are available in ABACUS when using NAOs as basis sets. These implementations differ primarily in their approach to projecting the electron density onto localized states, a crucial step for accurately modeling strong electron-electron interactions within specific atomic orbitals. The first approach is termed the “dual projection” method, while the second is referred to as the “full projection” method.
First, the dual projection method Qu et al. (2022a); Han, Ozaki, and Yu (2006) employs a Mulliken charge projector to construct the on-site density occupancy matrix. This approach facilitates the transformation of Kohn-Sham orbitals into a localized correlated subspace using NAOs, particularly for $d$ or $f$ orbitals. The Mulliken projector satisfies the charge sum rule, ensuring conservation of total electronic charge across all projected channels. This methodology is alternatively termed as the “dual orbitals” approach Han, Ozaki, and Yu (2006) due to its fundamental reliance on dual orbitals that maintain mutual orthogonality. Mathematically, these dual orbitals are derived from the original atomic orbitals. The spin-dependent on-site density occupancy matrix is explicitly defined as
$n^{I\sigma}_{mm^{\prime}}$ $=\frac{1}{4N_{k}}\sum_{\mathbf{k}\beta mm^{\prime}\mu}\Big{(}S^{I}_{\beta m,\mu}(\mathbf{k})\rho^{I\sigma}_{\mu,\beta m^{\prime}}(\mathbf{k})+\rho^{I\sigma}_{\beta m,\mu}(\mathbf{k})S^{I}_{\mu,\beta m^{\prime}}(\mathbf{k})$ (102)
$+S^{I}_{\beta m^{\prime},\mu}(\mathbf{k})\rho^{I\sigma}_{\mu,\beta m}(\mathbf{k})+\rho^{I\sigma}_{\beta m^{\prime},\mu}(\mathbf{k})S^{I}_{\mu,\beta m}(\mathbf{k})\Big{)},$
where $\rho^{I\sigma}_{\beta m,\mu}(\mathbf{k})$ denotes the density matrix, and $S^{I}_{\beta m,\mu}(\mathbf{k})$ is the overlap matrix of projected dual orbital $|\beta^{I}_{m}\rangle$ and numeric atomic orbital $|\phi_{\mu}\rangle$ (Eq. 83). More details are provided in Ref. Qu et al., 2022a.
Second, while the dual projection method offers a robust and efficient DFT+$U$ implementation, it relies on a subset of NAOs as projection operators introduces notable basis set dependence for the Hubbard $U$ parameter. This dependence becomes particularly evident when employing large radial cutoffs, which may affect the description of localized electron characteristics. Therefore, an alternative approach, implemented in ABACUS since version 3.6, involves the modulation of NAOs to construct localized projection operators. This method employs spherical truncation through empirical atomic radius settings, systematically transforming the problem into orbital modulation while satisfying three criteria. First, close correspondence with the original numerical orbitals must be ensured. Second, the orbitals should satisfy the normalization condition. Third, sufficient smoothness at the cutoff of the atomic radius must be ensured.
To be specific, the radial function $\chi(r)$ of the original NAOs undergoes direct truncation followed by normalization operations to yield the truncated localized orbital
@@
$\alpha(r)=\left.\frac{\chi(r)g(r;\gamma)}{\langle\chi(r)g(r;\gamma)|\chi(r)g(r;\gamma)\rangle}\right|_{\frac{\partial\langle\alpha|\chi\rangle}{\partial\gamma}=0},$ (103)
where the smooth function has the form
$g(r;\gamma)=\left\{\begin{matrix}1-\exp\left(-\frac{(r-r_{c})^{2}}{2\gamma^{2}}\right)&r<r_{c}\\ 0&r\geq{}r_{c}\end{matrix}\right.$. (104)
Here the parameter $\gamma$ controls the smoothing interval and the optimal $\gamma$ is determined iteratively to maximize two-center integrals between modulated $\alpha(r)$ and the original $\chi(r)$ orbitals, yielding the target localized orbitals. The full projection method constructs the on-site density occupancy matrix via the projection operator in real space Han, Ozaki, and Yu (2006)
@@
Figure 25: Band gaps of transition metal monoxide (a) MnO, (b) FeO, (c) CoO, and (d) NiO as influenced by the Hubbard $U$ parameter, utilizing a rhombohedral cell with a type-II antiferromagnetic structure. “ABACUS-Full” presents results from the full projection method. For comparative analysis, PBE+$U$ results for other computational methods are from Ref. Qu et al., 2022a.
We present systematic benchmarks of the full projection DFT+U method for MnO, CoO, FeO, and NiO systems using ABACUS v3.8. All calculations employ the optimized norm-conserving Vanderbilt (ONCV) pseudopotentials with SG15-v1.0 version. The Brillouin zone discretization is 0.15 $\text{bohr}^{-1}$ and the energy cutoff is 100 Ry. We use the PBE functional Perdew, Burke, and Ernzerhof (1996) with the Hubbard U correction, while spin-orbit coupling effects are excluded from consideration. In addition, we adopt a double-$\zeta$ plus polarization (DZP) atomic basis with a radial cutoff of 9.0 bohr for transition metal (TM) elements and 7.0 bohr for O atoms.
Table 1: Band gaps (in eV) and atomic magnetism (in $\mu_{B}$) with format “value of band gap (value of atomic magnetic of Mn/Co/Fe/Ni)” of MnO, FeO, CoO, and NiO as a function of effective on-site Coulomb energy $\bar{U}$ (in eV). The experimental values are presented in the last row.
@@
Exp. 3.63.6-$3.8$ Messick, Walker, and Glosser (1972) ($4.58$ Cheetham and Hope (1983)) $2.4$ Bowen, Adler, and Auker (1975) ($3.8$ Fender, Jacobson, and Wedgwood (1968)) $2.4$ Powell and Spicer (1970) ($3.32$ Roth (1958)) $4.0$ Sawatzky and Allen (1984)/$4.3$ Hüfner et al. (1984) ($1.90$ Khan and Erickson (1970))
Table 1 presents the calculated band gaps and atomic magnetic moments for transition metal oxides (TMOs) with the Hubbard U parameter ranging from 0 to 6 eV. All of the four investigated TMOs (MnO, FeO, CoO, and NiO) are in the rhombohedral structure with a type-II antiferromagnetic (AFM) configuration. The lattice parameters, adopted from Ref. Tran et al., 2006, are 4.445, 4.334, 4.254, and 4.171 Å for MnO, FeO, CoO, and NiO, respectively, with the AFM ordering aligned along the [111] crystallographic direction. Fig. 25 further shows the comparison of band gaps for MnO, FeO, CoO, and NiO systems. The results obtained from ABACUS match well with those from other DFT packages such as VASP, Wien2k, and Quantum ESPRESSO Qu et al. (2022a).
IV.7 Real-Time Time-Dependent DFT